OpenStudy (anonymous):

In how many ways can 3 cards be selected from a 12-card deck if the selected cards are not returned to the deck? I know the first step is 12C3 (c=combination) but what's my next step?

OpenStudy (anonymous):

compute \(_{12}C_3\) via \[\frac{12\times 11\times 10}{3\times 2}\]

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

I know that then the answer is 220, but I feel like there is more to the question

OpenStudy (anonymous):

What about the part of the question that says "if the selected cards are not returned to the deck"? Is that there just to trick me? @satellite73

OpenStudy (raden):

in selecting progress, one by one or direct 3 cards ?

OpenStudy (anonymous):

Selecting three at a time

OpenStudy (raden):

if a time,dont look returned or not :)

OpenStudy (anonymous):

"if the selected cards are not returned to the deck" just means that the total of cads will be less than before

OpenStudy (anonymous):

I understand that but after 12C3 then what do I do?

OpenStudy (anonymous):

@RadEn @Luis_Rivera @satellite73

OpenStudy (raden):

do u find the probability of that event ?

OpenStudy (anonymous):

That's irrelevant It's not asking me to find the probability

OpenStudy (raden):

well, it has cleared from u and satellite :)

OpenStudy (kropot72):

Your question is basically the same as "How many ways can three people be selected from a group of twelve people?" In general if the question states "selected" a combination is appropriate whereas if "arrangements" is stated a permutation applies. Therefore, as already answered, the solution to your question is 12C3. No other step is needed after that calculation